40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 15 01:17:10 2011, Keiko, Update, LSC, MICH locked and attempt to lock PRCL 
    Reply  Thu Sep 15 02:18:19 2011, Anamaria, Update, LSC, MICH locked and attempt to lock PRCL 
Message ID: 5413     Entry time: Thu Sep 15 01:17:10 2011     Reply to this: 5414
Author: Keiko 
Type: Update 
Category: LSC 
Subject: MICH locked and attempt to lock PRCL 

 Anamaria, Keiko

- We aligned MICH and were successfully locked MICH using AS55Q. The other mirrors were misaligned so that the other degrees of freedom didn't exist. AS55 was fed back to BS. The f2a filters on BS suspension were required to lock, because the pos feedback was unbalanced to angle degrees of freedom.

- We tried to lock PRCL next, however, because we aligned the MICH and the REFL beam paths were changed, REFL PDs didn't have the light anymore. The REFL paths were modified now, so we will try the PRCL locking next.

- We couldn't confirm REFL55 signals although we alined the REFL paths to REFL55 PD.

ELOG V3.1.3-